# Deep FlexQP: Accelerated Nonlinear Programming via Deep Unfolding

### v1

@@ Appendix G Further Details on QP Problem Classes
G.1 Random QPs
The first type of problems we study are random QPs of the form Equation 3. These are helpful as we can freely adjust the number of constraints as well as the sparsity of the problem directly in order to benchmark the optimizers under different operating conditions.
Problem Instances: We adopt the problem generation procedure from saravanos2025deep, where $P=M^{\top}M+\alpha I$ with $\alpha=1$ and all elements of $M$, $q$, $G$, and $A$ are standard normal distributed, i.e., each element $M_{ij},q_{i},G_{ij},A_{ij}\sim\mathcal{N}(0,1)$. The vectors $h$ and $b$ are generated using $h=G\xi$ and $b=A\zeta$ with $\xi,\zeta$ standard normal vectors.
@@ G.1 Random QPs
G.2 Portfolio Optimization
Portfolio Optimization is a foundational problem in finance where the goal is to maximize the risk-adjusted return of a group of assets (markowitz1952portfolio; boyd2013performance; boyd2017multi). This can be represented as the following QP (boyd2004convex; stellato2020osqp):
$\max_{x}\quad$ $\mu^{\top}x-\gamma(x^{\top}\Sigma x),$ (41a)
subject to $\mathbf{1}^{\top}x=1,$ (41b)
@@ G.2 Portfolio Optimization
$\min_{x,y}\quad$ $x^{\top}Dx+y^{\top}y-\gamma^{-1}\mu^{\top}x,$ (42a)
subject to $y=F^{\top}x,$ (42b)
@@
Problem Instances: For the medium-scale problems, we use the problem generation described in saravanos2025deep, setting $n=250$, $k=25$, and $\gamma=1.0$. The large-scale problems are generated using $n=10000$ assets, $k=100$ factors, and $\gamma=1.0$. The expected returns $\mu$ are sampled using $\mu_{i}\sim\mathcal{N}(0,1)$. The factor loading matrix $F$ has 50% non-zero elements sampled through $F_{ij}\sim\mathcal{N}(0,1)$. The diagonal elements of $D$ are generated uniformly as $D_{ii}\sim\mathcal{U}(0,\sqrt{k})$.
G.3 Support Vector Machines
Support Vector Machines (SVMs) is a classical machine learning problem where the goal is to find a linear classifier that best separates two sets of points (cortes1995support):
$\min_{x}x^{\top}x+\lambda\sum_{i=1}^{m}\max(0,b_{i}a_{i}^{\top}x+1),$ (43)
@@ G.3 Support Vector Machines
QP representation: The SVM problem Equation 43 can be converted into an equivalent QP representation (stellato2020osqp):
$\min_{x,t}\quad$ $x^{\top}x+\lambda\mathbf{1}^{\top}t,$ (44a)
subject to $t\geq\operatorname{diag}(b)Ax+1,$ (44b)
@@
This QP has $n+m$ decision variables and $2m$ inequalty constraints.
Problem Instances: We generate medium-scale problems using the rules from stellato2020osqp with $n=10$ features, $m=200$ data points, and $\lambda=1$. Large-scale problems are generated using $n=100$ features, $m=10000$ data points, and $\lambda=1$. The labels $b$ are chosen using
@@
G.4 LASSO
LASSO (least absolute shrinkage and selection operator) is a fundamental problem in statistics and machine learning (tibshirani1996regression; candes2008enhancing). The objective is to select sparse coefficients of a linear model that best match the given observations:
$\min_{x}\left\lVert Ax-b\right\rVert_{2}^{2}+\lambda\left\lVert x\right\rVert_{1},$ (47)
@@ G.4 LASSO
QP Representation: LASSO can be represented as a QP by introducing two extra decision variables $y\in\mathbb{R}^{m}$ and $t\in\mathbb{R}^{n}$ which help simplify the objective (stellato2020osqp):
$\min_{x,y,t}\quad$ $y^{\top}y+\lambda\mathbf{1}^{\top}t,$ (48a)
subject to $y=Ax-b,$ (48b)
@@
Problem Instances: We use the data generation procedure from (stellato2020osqp), where $A$ has 15% non-zero normally-distributed elements $A_{ij}\sim\mathcal{N}(0,1)$ and $b$ is generated through $b=Av+\epsilon$ with
@@
G.5 Huber Fitting
Huber Fitting is a robust least squares problem where the goal is to perform a linear regression with the assumption that outliers are present in the data (huber1964robust; huber1981robust):
$\min_{x}\sum_{i=1}^{m}\phi_{\text{hub}}(a_{i}^{\top}x-b_{i}),$ (50)
where the penalty function $\phi_{\text{hub}}$ penalizes the residuals quadratically when they are large and linearly when they are small:
@@ G.5 Huber Fitting
QP Representation: This robust least squares problem can be represented in the following QP form (stellato2020osqp):
$\min_{x,u,r,s}\quad$ $u^{\top}u+2\delta\mathbf{1}^{\top}(r+s),$ (52a)
subject to $Ax-b-u=r-s,$ (52b)
@@
Problem Instances: We follow stellato2020osqp and generate $A$ with 15% nonzero elements with $A_{ij}\sim\mathcal{N}(0,1)$ and set $b=Av+\epsilon$ where
@@
G.6 Linear Optimal Control
@@
$\min_{x,u}\quad$ $\sum_{t=0}^{T-1}x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}+x_{T}^{\top}Q_{T}x_{T},$ (54a)
subject to $x_{t+1}=A_{d}x_{t}+B_{d}u_{t},$ (54b)
@@
We study three classes of linear optimal control problems (OCPs). The first, Random Linear OCPs, consists of randomly generated stabilizable dynamics along with random costs, constraints, and initial conditions. The second and third classes, Double Integrator and Oscillating Masses, are adapted from chen2022large and contain dynamics with true physical interpretations. The randomness in these problems is given by sampling varying initial conditions for the systems as in saravanos2025deep.
G.6.1 Random Linear OCPs
We use the problem generation procedure similar to that in stellato2020osqp. We set the state dimension $n_{x}=8$ and $n_{u}=n_{x}/2=4$. The dynamics are generated by $A_{d}=X^{-1}AX$, where $A=\operatorname{diag}(a)\in{\mathbb{R}}^{n_{x}\times n_{x}}$ such that $a_{i}\sim{\mathcal{U}}(-1,1)$ and $X\in{\mathbb{R}}^{n_{x}\times n_{x}}$ with elements generated by $X_{ij}\sim{\mathcal{N}}(0,1)$, and $B_{d}\in{\mathbb{R}}^{n_{x}\times n_{u}}$ with $(B_{d})_{ij}\sim{\mathcal{N}}(0,1)$.
@@ G.6.1 Random Linear OCPs
G.6.2 Double Integrator
For the double integrator, adapted from chen2022large, we have $n_{x}=2$, $n_{u}=1$, and $T=20$ timesteps. The dynamics are fixed with
@@ G.6.3 Oscillating Masses
Appendix H Nonconvex Nonlinear Programming using SQP
H.1 Nonlinear Optimal Control
@@
$\begin{aligned} \operatorname*{minimize}_{x,u}\quad&\sum_{t=0}^{T-1}\ell(x_{t},u_{t})+\phi(x_{T}),\\ \text{subject to}\quad&x_{t+1}=F(x_{t},u_{t}),\quad\forall t=0,\ldots,T-1,\\ &x_{0}=\bar{x}_{0},\\ &h(x_{t})\leq 0,\quad\forall t=0,\ldots,T,\\ &g(u_{t})\leq 0,\quad\forall t=0,\ldots,T-1,\end{aligned}$ (62)
where $x_{t}\in{\mathbb{R}}^{n}$ and $u_{t}\in{\mathbb{R}}^{m}$ are the states and controls, respectively. The function $\ell:{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\to{\mathbb{R}}$ is the running cost and $\phi:{\mathbb{R}}^{n}\to{\mathbb{R}}$ is the terminal cost. The time horizon is $T>0$ and $\bar{x}_{0}\in{\mathbb{R}}^{n}$ is the initial condition. The problem formulation in Equation 62 includes state and control constraints represented by the functions $h(x_{t})$ and $g(u_{t})$. In the next two subsections, we provide the specific nonlinear optimal control examples for the cases of the Dubins vehicle and quadrotor.
H.1.1 Dubins Vehicle
The Dubins vehicle is a dynamics model with a state $x=(p_{x},p_{y},\theta)\in{\mathbb{R}}^{3}$, where $p_{x}$ and $p_{y}$ are the vehicle’s position in the Cartesian plane and $\theta$ is its orientation. We use the unicycle formulation of the continuous-time Dubins vehicle dynamics $\dot{x}=f(x,u)$ adapted from siciliano2009unicycle, where the control is given by $u=(v,\omega)\in{\mathbb{R}}^{2}$. Here, $v$ is the forward velocity of the vehicle and $\omega$ is the steering velocity of the vehicle.
@@ H.1.1 Dubins Vehicle
We formulate a nonlinear optimal control problem following Equation 62. We discretize the continuous-time dynamics using the Euler discretization $x_{t+1}=F(x_{t},u_{t})=x_{t}+f(x_{t},u_{t})\Delta t$ and use quadratic costs $\ell(x,u)=x^{\top}Qx+u^{\top}Ru$ and $\phi(x)=x^{\top}Q_{T}x$, where $Q=\operatorname{diag}(1.0,1.0,0.1)$, $R=0.1\cdot I$, and $Q_{T}=100\cdot Q$. The initial and target state, $x_{0}$ and $x_{\text{target}}$, are sampled uniformly from $\mathcal{U}(-\bar{x},\bar{x})$, where $\bar{x}=(5.0,5.0,\pi)$. The discretization of the dynamics uses $\Delta t=0.033$ and the time horizon for trajectory optimization is $T=50$ timesteps. We generate 5 circular obstacles with the form
$g_{i}(x_{t})=r_{i}^{2}-\left\lVert x_{t}-c_{i}\right\rVert_{2}^{2}\leq 0,\quad\forall t=0,\ldots,T,$
@@
[…]
H.1.2 Quadrotor
We use the continuous-time quadrotor dynamics model $\dot{x}=f(x,u)$ from Sabatino2015QuadrotorCM. The model consists of the state $x\in{\mathbb{R}}^{12}$ that includes the linear positions, angles, linear velocities, and angular velocities. The system is actuated by a four controls, the collective thrust $F$ and three torques, given by $u=(F,\tau_{x},\tau_{y},\tau_{z})\in{\mathbb{R}}^{3}$.
@@ H.1.2 Quadrotor
Using this model, we formulate nonlinear optimal control problems as in Equation 62. We discretize the dynamics through the Euler discretization $x_{t+1}=F(x_{t},u_{t})=x_{t}+f(x_{t},u_{t})\Delta t$. The cost in Equation 62 is defined by the quadratic cost $\ell(x,u)=x^{\top}Qx+u^{\top}Ru$ and $\phi(x)=x^{\top}Q_{T}x$, where the cost matrices are given by
@@
[…]
H.2 Nonlinear Predictive Safety Filters
@@
Finally, we apply our proposed approach to accelerate a predictive safety filter for nonlinear model predictive control. These methods are based on control barrier functions (CBFs) (ames2019control) and filter a reference control $u^{\text{ref}}$ so that it better respects safety constraints (wabersich2021predictive). The following optimization is solved at every MPC step:
@@
$\begin{aligned} \operatorname*{minimize}_{x,u}\quad&\sum_{t=0}^{T-1}\left\lVert u_{t}-u_{t}^{\text{ref}}\right\rVert_{2}^{2},\\ \text{subject to}\quad&x_{t+1}=F(x_{t},u_{t}),\quad\forall t=0,\ldots,T-1,\\ &x_{0}=\bar{x}_{0}\\ &(1-\beta)h(x_{t})-h(x_{t+1})\leq 0,\quad\forall t=0,\ldots,T-1,\\ &g(u_{t})\leq 0,\quad\forall t=0,\ldots,T-1,\end{aligned}$ (63)
@@
where $\beta\in(0,1)$ is a parameter controlling the strength of the CBF constraint. This differs from Equation 62 because the discrete CBF constraint is defined between two consecutive states $x_{t}$ and $x_{t+1}$ rather than assuming the state constraints are separable across time. Our method improves upon the Shield-MPPI method proposed by yin2023shield because our optimization explicitly incorporates the dynamics and input constraints while also minimizing the discrepancy from the reference control trajectory. Furthermore, using our accelerated Deep FlexQP ensures that the optimization can be run fast enough for real-time control. We use a version of Deep FlexQP with performance guarantees from minimizing the generalization bound loss (see Appendix L).
H.2.1 Shield-MPPI

### v2

@@ Appendix H Further Details on QP Problem Classes
H.1 Random QPs
The first type of problems we study are random QPs of the form Eq. 3. These are helpful as we can freely adjust the number of constraints as well as the sparsity of the problem directly in order to benchmark the optimizers under different operating conditions.
Problem Instances: We adopt the problem generation procedure from Saravanos et al. (2025), where $P=M^{\top}M+\alpha I$ with $\alpha=1$ and all elements of $M$, $q$, $G$, and $A$ are standard normal distributed, i.e., each element $M_{ij},q_{i},G_{ij},A_{ij}\sim\mathcal{N}(0,1)$. The vectors $h$ and $b$ are generated using $h=G\xi$ and $b=A\zeta$ with $\xi,\zeta$ standard normal vectors.
@@ H.1 Random QPs
H.2 Portfolio Optimization
Portfolio Optimization is a foundational problem in finance where the goal is to maximize the risk-adjusted return of a group of assets (Markowitz, 1952; Boyd et al., 2013; 2017). This can be represented as the following QP (Boyd and Vandenberghe, 2004; Stellato et al., 2020):
$\operatorname*{maximize}_{x}\quad$ $\mu^{\top}x-\gamma(x^{\top}\Sigma x),$ (33a)
$\operatorname*{subject\ to}\quad$ $\mathbf{1}^{\top}x=1,$ (33b)
@@ H.2 Portfolio Optimization
$\operatorname*{minimize}_{x,y}\quad$ $x^{\top}Dx+y^{\top}y-\gamma^{-1}\mu^{\top}x,$ (34a)
$\operatorname*{subject\ to}\quad$ $y=F^{\top}x,$ (34b)
@@
Problem Instances: For the medium-scale problems, we use the problem generation described in Saravanos et al. (2025), setting $n=250$, $k=25$, and $\gamma=1.0$. The large-scale problems are generated using $n=10000$ assets, $k=100$ factors, and $\gamma=1.0$. The expected returns $\mu$ are sampled using $\mu_{i}\sim\mathcal{N}(0,1)$. The factor loading matrix $F$ has 50% non-zero elements sampled through $F_{ij}\sim\mathcal{N}(0,1)$. The diagonal elements of $D$ are generated uniformly as $D_{ii}\sim\mathcal{U}(0,\sqrt{k})$.
H.3 Support Vector Machines
Support Vector Machines (SVMs) are a classical machine learning method where the goal is to find a linear classifier that best separates two sets of points (Cortes and Vapnik, 1995):
$\operatorname*{minimize}_{x}x^{\top}x+\lambda\sum_{i=1}^{m}\max(0,b_{i}a_{i}^{\top}x+1),$ (35)
@@ H.3 Support Vector Machines
QP representation: The SVM problem Eq. 35 can be converted into an equivalent QP representation (Stellato et al., 2020):
$\operatorname*{minimize}_{x,t}\quad$ $x^{\top}x+\lambda\mathbf{1}^{\top}t,$ (36a)
$\operatorname*{subject\ to}\quad$ $t\geq\operatorname{diag}(b)Ax+1,$ (36b)
@@
This QP has $n+m$ decision variables and $2m$ inequality constraints.
Problem Instances: We generate medium-scale problems using the rules from Stellato et al. (2020) with $n=10$ features, $m=200$ data points, and $\lambda=1$. Large-scale problems are generated using $n=100$ features, $m=10000$ data points, and $\lambda=1$. The labels $b$ are chosen using
@@
H.4 LASSO
LASSO (least absolute shrinkage and selection operator) is a fundamental problem in statistics and machine learning (Tibshirani, 1996; Candes et al., 2008). The objective is to select sparse coefficients of a linear model that best match the given observations:
$\operatorname*{minimize}_{x}\left\lVert Ax-b\right\rVert_{2}^{2}+\lambda\left\lVert x\right\rVert_{1},$ (39)
@@ H.4 LASSO
QP Representation: LASSO can be represented as a QP by introducing two extra decision variables $y\in\mathbb{R}^{m}$ and $t\in\mathbb{R}^{n}$ which help simplify the objective (Stellato et al., 2020):
$\operatorname*{minimize}_{x,y,t}\quad$ $y^{\top}y+\lambda\mathbf{1}^{\top}t,$ (40a)
$\operatorname*{subject\ to}\quad$ $y=Ax-b,$ (40b)
@@
Problem Instances: We use the data generation procedure from (Stellato et al., 2020), where $A$ has 15% non-zero normally-distributed elements $A_{ij}\sim\mathcal{N}(0,1)$ and $b$ is generated through $b=Av+\epsilon$ with
@@
H.5 Huber Fitting
Huber Fitting is a robust least squares problem where the goal is to perform a linear regression with the assumption that outliers are present in the data (Huber, 1964; 1981):
$\operatorname*{minimize}_{x}\sum_{i=1}^{m}\phi_{\text{hub}}(a_{i}^{\top}x-b_{i}),$ (42)
where the penalty function $\phi_{\text{hub}}$ penalizes the residuals linearly when they are large and quadratically when they are small:
@@ H.5 Huber Fitting
QP Representation: This robust least squares problem can be represented in the following QP form (Stellato et al., 2020):
$\operatorname*{minimize}_{x,u,r,s}\quad$ $u^{\top}u+2\delta\mathbf{1}^{\top}(r+s),$ (44a)
$\operatorname*{subject\ to}\quad$ $Ax-b-u=r-s,$ (44b)
@@
Problem Instances: We follow Stellato et al. (2020) and generate $A$ with 15% nonzero elements with $A_{ij}\sim\mathcal{N}(0,1)$ and set $b=Av+\epsilon$ where
@@
H.6 Linear Optimal Control
@@
$\operatorname*{minimize}_{x,u}\quad$ $\sum_{t=0}^{T-1}x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t}+x_{T}^{\top}Q_{T}x_{T},$ (46a)
$\operatorname*{subject\ to}\quad$ $x_{t+1}=A_{d}x_{t}+B_{d}u_{t},$ (46b)
@@
We study three classes of linear optimal control problems (OCPs). The first, Random Linear OCPs, consists of randomly generated stabilizable dynamics along with random costs, constraints, and initial conditions. The second and third classes, Double Integrator and Oscillating Masses, are adapted from Chen et al. (2022a) and contain dynamics with true physical interpretations. The randomness in these problems is given by sampling varying initial conditions for the systems as in Saravanos et al. (2025).
H.6.1 Random Linear OCPs
We use the problem generation procedure similar to that in Stellato et al. (2020). We set the state dimension $n_{x}=8$ and $n_{u}=n_{x}/2=4$. The dynamics are generated by $A_{d}=X^{-1}AX$, where $A=\operatorname{diag}(a)\in{\mathbb{R}}^{n_{x}\times n_{x}}$ such that $a_{i}\sim{\mathcal{U}}(-1,1)$ and $X\in{\mathbb{R}}^{n_{x}\times n_{x}}$ with elements generated by $X_{ij}\sim{\mathcal{N}}(0,1)$, and $B_{d}\in{\mathbb{R}}^{n_{x}\times n_{u}}$ with $(B_{d})_{ij}\sim{\mathcal{N}}(0,1)$.
@@ H.6.1 Random Linear OCPs
H.6.2 Double Integrator
For the double integrator, adapted from Chen et al. (2022a), we have $n_{x}=2$, $n_{u}=1$, and $T=20$ timesteps. The dynamics are fixed with
@@ H.6.3 Oscillating Masses
Appendix I Nonconvex Nonlinear Programming using SQP
I.1 Nonlinear Optimal Control
@@
$\operatorname*{minimize}_{x,u}\quad$ $\sum_{t=0}^{T-1}\ell(x_{t},u_{t})+\phi(x_{T}),$ (54a)
$\operatorname*{subject\ to}\quad$ $x_{t+1}=F(x_{t},u_{t}),\quad\forall t=0,\ldots,T-1,$ (54b)
$x_{0}=\bar{x}_{0},$ (54c)
$g(x_{t})\leq 0,\quad\forall t=0,\ldots,T,$ (54d)
$h(u_{t})\leq 0,\quad\forall t=0,\ldots,T-1,$ (54e)
where $x_{t}\in{\mathbb{R}}^{n}$ and $u_{t}\in{\mathbb{R}}^{m}$ are the states and controls, respectively. The function $\ell:{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\to{\mathbb{R}}$ is the running cost and $\phi:{\mathbb{R}}^{n}\to{\mathbb{R}}$ is the terminal cost. The time horizon is $T>0$ and $\bar{x}_{0}\in{\mathbb{R}}^{n}$ is the initial condition. The problem formulation in Eq. 54 includes state and control constraints represented by the functions $g(x_{t})$ and $h(u_{t})$. In the next two subsections, we provide the specific nonlinear optimal control examples for the cases of the Dubins vehicle and quadrotor.
I.1.1 Dubins Vehicle
The Dubins vehicle is a dynamics model with a state $x=(p_{x},p_{y},\theta)\in{\mathbb{R}}^{3}$, where $p_{x}$ and $p_{y}$ are the vehicle’s position in the Cartesian plane and $\theta$ is its orientation. We use the unicycle formulation of the continuous-time Dubins vehicle dynamics $\dot{x}=f(x,u)$ adapted from Siciliano et al. (2009), where the control is given by $u=(v,\omega)\in{\mathbb{R}}^{2}$. Here, $v$ is the forward velocity of the vehicle and $\omega$ is the steering velocity of the vehicle.
@@ I.1.1 Dubins Vehicle
We formulate a nonlinear optimal control problem following Eq. 54. We discretize the continuous-time dynamics using the Euler discretization $x_{t+1}=F(x_{t},u_{t})=x_{t}+f(x_{t},u_{t})\Delta t$ and use quadratic costs $\ell(x,u)=x^{\top}Qx+u^{\top}Ru$ and $\phi(x)=x^{\top}Q_{T}x$, where $Q=\operatorname{diag}(1.0,1.0,0.1)$, $R=0.1\cdot I$, and $Q_{T}=100\cdot Q$. The initial and target state, $x_{0}$ and $x_{\text{target}}$, are sampled uniformly from $\mathcal{U}(-\bar{x},\bar{x})$, where $\bar{x}=(5.0,5.0,\pi)$. The discretization of the dynamics uses $\Delta t=0.033$ and the time horizon for trajectory optimization is $T=50$ timesteps. We generate 5 circular obstacles with the form
$g_{i}(x_{t})=r_{i}^{2}-\left\lVert x_{t}-c_{i}\right\rVert_{2}^{2}\leq 0,\quad\forall t=0,\ldots,T,$ (55)
@@
For generating the QP training data, we generate 500 QP subproblems by solving randomly generated Dubins vehicle problems with SQP using OSQP as the QP solver. For evaluation, we generate 100 random control problems and solve them using SQP with OSQP or SQP with Deep FlexQP. Each algorithm is allowed 50 SQP iterations and runs until the infinity norm of the SQP residuals falls below an absolute tolerance of $\varepsilon=10^{-2}$. Furthermore, each QP solver runs until convergence of $10^{-3}$ is reached, with a max budget of 10 seconds and an unlimited number of iterations.
I.1.2 Quadrotor
We use the continuous-time quadrotor dynamics model $\dot{x}=f(x,u)$ from Sabatino (2015). The model consists of the state $x\in{\mathbb{R}}^{12}$ that includes the linear positions, angles, linear velocities, and angular velocities. The system is actuated by four controls, the collective thrust $F$ and three torques, given by $u=(F,\tau_{x},\tau_{y},\tau_{z})\in{\mathbb{R}}^{4}$.
@@ I.1.2 Quadrotor
Using this model, we formulate nonlinear optimal control problems as in Eq. 54. We discretize the dynamics through the Euler discretization $x_{t+1}=F(x_{t},u_{t})=x_{t}+f(x_{t},u_{t})\Delta t$. The cost in Eq. 54 is defined by the quadratic cost $\ell(x,u)=x^{\top}Qx+u^{\top}Ru$ and $\phi(x)=x^{\top}Q_{T}x$, where the cost matrices are given by
@@
For generating the QP training data, we generate 500 QP subproblems by solving randomly generated quadrotor problems with SQP using OSQP as the QP solver. For evaluation, we generate 100 random quadrotor problems and solve them using SQP with OSQP or SQP with Deep FlexQP. Similar to the Dubins vehicle, each algorithm is allowed 50 SQP iterations and success in Fig. 7 (left) is achieved when the infinity norm of the SQP residuals falls below the absolute convergence tolerance $\varepsilon=10^{-2}$. Each QP solver runs until convergence of $10^{-3}$ is reached, with a max budget of 10 seconds and an unlimited number of iterations.
I.2 Nonlinear Predictive Safety Filters
@@
Finally, we apply our proposed approach to accelerate a predictive safety filter for nonlinear model predictive control. These methods are based on control barrier functions (CBFs) (Ames et al., 2019) and filter a reference control $u^{\text{ref}}$ so that it better respects safety constraints (Wabersich and Zeilinger, 2021). The following optimization is solved at every MPC step:
@@
$\operatorname*{minimize}_{x,u}\quad$ $\sum_{t=0}^{T-1}\left\lVert u_{t}-u_{t}^{\text{ref}}\right\rVert_{2}^{2},$ (56a)
$\operatorname*{subject\ to}\quad$ $x_{t+1}=F(x_{t},u_{t}),\quad\forall t=0,\ldots,T-1,$ (56b)
$x_{0}=\bar{x}_{0}$ (56c)
$(1-\beta)g(x_{t})-g(x_{t+1})\leq 0,\quad\forall t=0,\ldots,T-1,$ (56d)
$h(u_{t})\leq 0,\quad\forall t=0,\ldots,T-1,$ (56e)
@@
where $\beta\in(0,1)$ is a parameter controlling the strength of the CBF constraint. This differs from Eq. 54 because the discrete CBF constraint is defined between two consecutive states $x_{t}$ and $x_{t+1}$ rather than assuming the state constraints are separable across time. Our method improves upon the Shield-MPPI method proposed by Yin et al. (2023) because our optimization explicitly incorporates the dynamics and input constraints while also minimizing the discrepancy from the reference control trajectory. Furthermore, using our accelerated Deep FlexQP ensures that the optimization can be run fast enough for real-time control. We use a version of Deep FlexQP with performance guarantees from minimizing the generalization bound loss (see Appendix M).
I.2.1 Shield-MPPI